Circular No. 7652

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

THE EDGAR WILSON AWARD With reference to IAUC 6936 (and also to IAUC 7223 and 7445), the

Smithsonian Astrophysical Observatory wishes to announce that the 2001 Edgar Wilson Award for the discovery of comets is being divided among the following two amateur astronomers for C/2000 W1: Albert Francis Arthur Lofley Jones, Stoke, Nelson, New Zealand; and Syogo Utsunomiya, Minami-Oguni machi, Aso-gun, Kumamoto-ken, Japan.

SUPERNOVA 2001cx IN UGC 12266 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum of SN 2001cx (cf. IAUC 7651), obtained by M. Calkins on June 28.44 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova past maximum light. The spectrum exhibits a flat continuum with well-defined P-Cyg lines of hydrogen, consistent with the plateau phase of a normal type-II supernova. Adopting the NASA/IPAC Extragalactic Database recession velocity of 4817 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is ∼ 4800 km/s.

SUPERNOVA 2001cp IN UGC 10738 M. Dennefeld, Institut d’Astrophysique de Paris, writes: “A calibrated

low-dispersion spectrum (range 370–733 nm; resolution 0.6 nm) of SN 2001cp (cf. IAUC 7645, 7649) was obtained with the 1.93-m telescope at Haute-Provence Observatory on June 21.97 UT. The spectrum shows a blue continuum, with strong Ca II/Si II absorption and emission features in the blue — along with a marked, but comparatively more shallow, Si II 635-nm absorption, indicating a type-Ia supernova reaching maximum. The spectrum in fact closely resembles that of SN 1990N a week before maximum. With a recession velocity of the parent galaxy of 6719 km/s taken from the Lyon Extragalactic Database, the Si II 635-nm absorption yields an expansion velocity of 11 000 km/s, while the blue Ca II lines give 13 000–16 000 km/s. Another spectrograph obtained on June 24.98 shows no significant spectral change since June 21, with the exception of a flattening of the continuum slope.”

G. Masi, Ceccano, Italy, reports an unfiltered magnitude of 16.1 from a CCD image obtained on June 22.937 UT. He also provides the following accurate position for SN 2001cp: α = 17^{h}11^{m}02^{s}.65, δ = +5^{o}50′27′′.8 (equinox 2000.0; eight USNO-A2.0 reference stars).

2001 June 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green